i dont get how to do this at all. i learned once, but ive forgotten & i forgot to bring home my book. the problem is what is the solution of the following system of linear equations?
y = 2x
4x - y = 10
can you please explain to me how to do this, so i can do it and the rest of them? :]

Answer by stanbon(75887)

y = 2x
4x - y = 10

Sinc y=2x, substitute that value into the 2nd equation to get:
4x-(2x)=10
2x=10
x=5
Substitute that into the 1st equation to get:
y=2(5)=10
Solution: x=5, y=10
Cheers,
Stan H.